The pressure inside a container is 77 atm at a temperature of 570C. What would the pressure be at 75o C?


1
Expert's answer
2021-05-25T07:51:35-0400

"P_1\/T_1=P_2\/T_2"

"P_1=77atm, T_2=57+273=330K"

"P_2=?, T_2=75+273=348K"

"77\/330=P_2\/348"

"P_2=\frac{77\times 348}{330}=81.2atm"